The treasurer of Tropical Fruits, Inc., has projected the cash flows of Projects A, B, and C as follows:

Year Project A Project B Project C
0 −$ 215,000 −$ 380,000 −$ 215,000
1 138,000 232,000 152,000
2 138,000 232,000 118,000


Suppose the relevant discount rate is 9 percent per year.

a. Compute the profitability index for each of the three projects. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Profitability
index
Project A
Project B
Project C


b. Compute the NPV for each of the three projects. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

NPV
Project A $
Project B $
Project C $


c. Suppose these three projects are independent. Which project(s) should the company accept based on the profitability index rule?

  • Project A

  • Project B

  • Project C

  • Project A, Project B, Project C

  • Project A, Project B

  • Project A, Project C

  • Project B, Project C


d. Suppose these three projects are mutually exclusive. Which project(s) should the company accept based on the profitability index rule?

  • Project A

  • Project B

  • Project C

  • Project A, Project B, Project C

  • Project A, Project B

  • Project A, Project C

  • Project B, Project C


e. Suppose the budget for these projects is $595,000. The projects are not divisible. Which project(s) should be accepted?

  • Project A

  • Project B

  • Project C

  • Project A, Project B, Project C

  • Project B, Project C

  • Project B, Project A

  • Project A, Project C
Answer #1

a. Compute the profitability index for each of the three projects

-Project A

Year Cashflow PVF@9% Cashflow*PVF
0           (215,000) 1            (215,000.00)
1              138,000 0.9174              126,605.50
2              138,000 0.8417              116,151.84

Profitability index = PV of future Cash Flows / Initial Investment

= (126605.50+116151.84)/215000

= 242757.34/215000

= 1.13

-Project B

Year Cashflow PVF@9% Cashflow*PVF
0           (380,000) 1            (380,000.00)
1              232,000 0.9174              212,844.04
2              232,000 0.8417              195,269.76

Profitability index = PV of future Cash Flows / Initial Investment

= (212844.04+195269.76)/380000

= 408113.80/380000

= 1.07

-Project C

Year Cashflow PVF@9% Cashflow*PVF
0           (215,000) 1            (215,000.00)
1              152,000 0.9174              139,449.54
2              118,000 0.8417                99,318.24

Profitability index = PV of future Cash Flows / Initial Investment

= (139449.54+99318.24)/215000

= 238767.78/215000

= 1.11

b. Compute the NPV for each of the three projects.

- Project A

NPV = PV of Inflows - PV of Outflows

= (126605.50+116151.84)-215000

= 242757.34-215000

= 27757.34

- Project B

NPV = PV of Inflows - PV of Outflows

= (212844.04+195269.76)-380000

= 408113.80-380000

= 28113.80

​​​​​​​- Project C

NPV = PV of Inflows - PV of Outflows

= (139449.54+99318.24)-215000

= 238767.78-215000

= 23767.78

c. Suppose these three projects are independent. Which project(s) should the company accept based on the profitability index rule?

Since the projects are independent, we can select all projects having profitability index greater than 1, provided the availability of enough fund. Here all the three projects have profitability index greater than 1. So accept Project A, Project B, Project C​​​​​​​.

d. Suppose these three projects are mutually exclusive. Which project(s) should the company accept based on the profitability index rule?

Since the projects are mutually exclusive, select the project with higher profitability index. So accept Project A.

e. Suppose the budget for these projects is $595,000. The projects are not divisible. Which project(s) should be accepted?

Select those having higher NPV. Project B, Project A